# Shadows of the Colossus: Hierarchical Black Hole Mergers in a 10-million-body Globular Cluster Simulation

### v1

@@ 5.3 Predicted Virgo cluster BBH merger populations
We utilize Equation 10 to calculate $N_{\rm bbh}$ and $N_{\rm 2G+}$ for each GC in the Virgo cluster. Figure 11 again plots the mass-$r_{h}$ parameter space of Virgo GCs (excluding six GCs with mass exceeding the initial mass of colossus), colored according to the predicted number of total BBH mergers ($N_{\rm bbh}$, left panel) and the predicted number of generation 2+ hierarchical mergers ($N_{\rm 2G+}$, right panel). The total number of mergers in each panel is shown to increase as the cluster density increases, as expected. For very low-mass clusters with present-day masses less than roughly $10^{4.5}\,M_{\odot}$, the predicted number of mergers can be seen to increase slightly. This is because some of these clusters are best matched by CMC models that were initially more massive (and therefore formed many BBH mergers, especially at early times), but lost a significant amount of mass via enhanced tidal stripping in their galactic potential (see Kremer et al., 2020b, for further discussion).
@@ 6.2 Future Work
We have focused here on black hole growth via hierarchical mergers, however other growth channels are likely possible in dense stellar clusters. One scenario is growth via massive stellar mergers and collisions while the cluster is very young ($t\lesssim 10\,$Myr). A number of studies have demonstrated these may provide a pathway for forming black holes in or beyond the pair-instability mass gap (e.g., Di Carlo et al., 2019; Kremer et al., 2020c; González et al., 2021; Costa et al., 2022; Ballone et al., 2023). Indeed, $767$ black holes with masses in the pair-instability gap formed via binary stellar collisions/mergers in colossus (see the small handful of G1 points in Figure 3 with primary mass $40.5\,M_{\odot}$ which formed via this channel). Additionally, several studies have shown that black holes may also grow via accretion upon collision with other massive stars (e.g., Giersz et al., 2015; Kıroğlu et al., 2025a, b). Rose et al. (2023) pointed out that this channel may be especially prominent in very dense star clusters, with densities similar to the Galactic center. In colossus, we identify $759$ black hole+star collisions events over the cluster’s lifetime. For the vast majority of these ($90\%$), the $M_{\star}/M_{\rm bh}$ mass ratio is $0.1$ or less. In this regime, significant black hole growth is unlikely (Kremer et al., 2022). Only five collisions feature $M_{\star}/M_{\rm bh}$ values of 0.5 or higher (all occurring before 2525\,Myr) where significant growth may be feasible. However, we have assumed in colossus a $5\%$ binary fraction for all stars. Realistic clusters may very well have much higher binary fractions, especially for massive stars (e.g., Offner et al., 2023). Higher binary fractions may significantly increase the number of massive star–black hole collisions (Kıroğlu et al., 2025a). We reserve exploration of this possibility for future models.
@@
The authors thank Carl Rodriguez and Fred Rasio for helpful discussions throughout the preparation of the manuscript. A.M. acknowledges support from the Undergraduate Research Scholarship (URS) summer program at UCSD. F.K. acknowledges support from the CIERA Postdoctoral Fellowship.

### v2

@@ 5.3 Predicted Virgo cluster BBH merger populations
We utilize Equation 10 to calculate $N_{\rm bbh}$ and $N_{\rm 2G+}$ for each GC in the Virgo cluster. Figure 11 again plots the mass-$r_{h}$ parameter space of Virgo GCs (excluding six GCs with mass exceeding the initial mass of colossus), colored according to the predicted number of total BBH mergers ($N_{\rm bbh}$, left panel) and the predicted number of generation 2+ hierarchical mergers ($N_{\rm 2G+}$, right panel). The total number of mergers in each panel is shown to increase as the cluster density increases, as expected. For very low-mass clusters with present-day masses less than roughly $10^{4.5}\,M_{\odot}$, the predicted number of mergers can be seen to increase slightly. This is because some of these clusters are best matched by CMC models that were initially more massive (and therefore formed many BBH mergers, especially at early times), but lost a significant amount of mass via enhanced tidal stripping in their galactic potential (see kremer2020modeling, for further discussion).
@@ 6.2 Future Work
We have focused here on black hole growth via hierarchical mergers, however other growth channels are likely possible in dense stellar clusters. One scenario is growth via massive stellar mergers and collisions while the cluster is very young ($t\lesssim 10\,$Myr). A number of studies have demonstrated these may provide a pathway for forming black holes in or beyond the pair-instability mass gap (e.g., DiCarlo2019; Kremer2020; Gonzalez2021; Costa2022; Ballone2023). Indeed, $767$ black holes with masses in the pair-instability gap formed via binary stellar collisions/mergers in colossus (see the small handful of G1 points in Figure 3 with primary mass $40.5\,M_{\odot}$ which formed via this channel). Additionally, several studies have shown that black holes may also grow via accretion upon collision with other massive stars (e.g., Giersz2015; Kiroglu2025a; Kiroglu2025c). Rose2023 pointed out that this channel may be especially prominent in very dense star clusters, with densities similar to the Galactic center. In colossus, we identify $759$ black hole+star collisions events over the cluster’s lifetime. For the vast majority of these ($90\%$), the $M_{\star}/M_{\rm bh}$ mass ratio is $0.1$ or less. In this regime, significant black hole growth is unlikely (Kremer2022_sph). Only five collisions feature $M_{\star}/M_{\rm bh}$ values of 0.5 or higher (all occurring before 2525\,Myr) where significant growth may be feasible. However, we have assumed in colossus a $5\%$ binary fraction for all stars. Realistic clusters may very well have much higher binary fractions, especially for massive stars (e.g., Offner2023). Higher binary fractions may significantly increase the number of massive star–black hole collisions (Kiroglu2025a). We reserve exploration of this possibility for future models.
@@
The authors thank Carl Rodriguez and Fred Rasio for helpful discussions throughout the preparation of the manuscript. We also thank the anonymous referee for helpful feedback. A.M. acknowledges support from the Undergraduate Research Scholarship (URS) summer program at UCSD. F.K. acknowledges support from the CIERA Postdoctoral Fellowship.